# GeSubNet: Gene Interaction Inference for Disease Subtype Network Generation

### v1

@@ Appendix D Hyperparameter Setting
Appendix E Evaluation Metrics
@@
(4) Number of enriched biological functions (#EBF). Corresponding to the differences in graph properties of gene networks, we also explore their biological significance. A commonly used method for this is functional enrichment analysis, which identifies biological functions, pathways, and molecular activities that are overrepresented within a gene set when compared to a random selection of genes with a similar size and degree distribution from the genome. In our study, we performed Gene Ontology (GO) enrichment analysis using the R package clusterProfiler (Ashburner et al., 2000), which leverages data from databases such as KEGG and GO to identify enriched biological terms. A greater degree of enrichment suggests that the network exhibits more meaningful gene interactions than would be expected by chance. This unique enrichment across subtypes implies that the gene networks represent biologically significant interactions, where genes within specific cancer subtype networks are functionally connected as a group. To evaluate the functional diversity between two gene networks, we conducted an experiment using GO to count the number of unique GO terms associated with the genes in each network. Specifically, we used the enrichGO() function from clusterProfiler to map the genes from both networks to their corresponding GO terms. The compareCluster() function was applied to compare the sets of GO terms associated with each network and to identify differences, focusing on the number of enriched biological functions. To quantify the differences, we calculated the number of enriched biological functions (#EBF) using the symmetric difference between the sets of GO terms. Mathematically, this is represented as: #$\#EBF=(GO(\mathcal{G}_{1})\setminus GO(\mathcal{G}_{2}))\cup(GO(\mathcal{G}_{2})\setminus GO(\mathcal{G}_{1}))$. This operation captures the unique functions present in one network but not another. Enrichment was evaluated based on statistical significance, where the biological functions with a p-value $<0.05$ were reported. A higher #EBF indicates that the networks capture different biological processes or molecular functions, potentially reflecting the underlying biological differences between the networks’ contexts.
Appendix F GO Function Enrichment Analysis
@@
Table 8 presents the enriched Gene Ontology (GO) terms associated with various biological functions across four cancer types (BRCA, GBM, LGG, and OV), as identified using different methods in the GO analysis of Experiment II. The Venn diagrams in Figure 3 illustrate the overlaps among the results from the different methods. Due to the complexity of comparing multiple methods, we present a four-way Venn diagram focusing on four selected methods (WGCNA, CSGNN, LR-GNN, and GenSubNet) for clarity.
GenSubNet findings: GenSubNet shows a robust array of enriched GO terms across different cancers, including:
@@
Comparison: The proposed method identifies a broader range of distinct GO terms compared to other methods, and the GO term set identified by GenSubNet constitutes a superset of the terms determined by different methods.
@@
Appendix G Simulated Gene Knockout Experiment
G.1 Workflow
@@ G.2 Shift Rate
Appendix H Case Study
H.1 Breast Invasive Carcinoma
@@
H.2 Experiments and Analysis Protocols
@@
Appendix I Ablation Studies
@@
Ablation. We conduct three detailed ablation studies to evaluate the impact of each module in GenSubNet. More details can be found in Appendix I. Figure 7 presents the results of the three ablation studies across all variant models. For Patient-M, the proposed sample encoder significantly outperforms all other DGM models across the four network inference tasks (BRCA, GBM, LGG, and OV). For instance, the proposed method achieves an average improvement of 32.3%/31.2%/22.1%/32.3% in terms of GED. The Graph-M ablations show that the method using Neo-GNN consistently performs best, while the other GNN models yield comparable results. For Infer-M ablation, GenSubNet significantly outperforms the other objective functions, achieving approximately twice the metric values of its counterparts.
@@
Appendix J Prior Graph V.S. Newly Generated Graph
We evaluated the performance of patient group learning by inputting the newly generated graph from the GenSubNet into a plain GCN and comparing the results. Figure 8 presents a UMAP visualization of the learned latent sample spaces, with the prior graph initialization (Left) and the generated graph GCN initialization (Right). The left sub-figure shows that different patient groups appear mixed in the latent sample space derived from the prior gene network. However, there are clearer boundaries between various patient groups, as shown on the right side. Such results confirm the redundancy of information in the common prior gene networks. It demonstrates that the GenSubNet provides more structured information and potential for cancer studies.

### v3

@@ Appendix D Hyperparameter Setting
Appendix E Computational Requirements
Table 8: The computational requirements of the proposed method, including both runtime (training and inference time) and memory usage, across different cancer datasets.
Dataset Training Time (sec) Inference Time (sec) GPU Memory Usage (MB)
BRCA 6.23 $\pm$ 0.03 2.21 $\pm$ 0.05 4393 $\pm$ 152
GBM 3.64 $\pm$ 0.04 1.44 $\pm$ 0.07 2764 $\pm$ 117
LGG 5.96 $\pm$ 0.03 1.98 $\pm$ 0.04 3834 $\pm$ 122
OV 3.45 $\pm$ 0.04 1.25 $\pm$ 0.04 2583 $\pm$ 108
Pan-cancer 454.47 $\pm$ 12.2 142.86 $\pm$ 5.2 15893 $\pm$ 733
We evaluated the computational requirements of the proposed method, including runtime (training and inference time) and memory usage, across four cancer datasets. While cancer patient gene expression datasets typically contain hundreds of patients for each cancer type, we assessed the method’s scalability by conducting an additional test using the TCGA Pan-Cancer expression dataset based on the BRCA graph. This dataset includes 8,314 patient samples and is one of the largest pan-cancer datasets. The results are summarized in the Table 8. We find that the computational requirements of the proposed method are manageable for real-world cancer datasets and scale effectively to larger datasets with more patient samples.
Appendix F Evaluation Metrics
@@
(4) Number of enriched biological functions (#EBF). Corresponding to the differences in graph properties of gene networks, we also explore their biological significance. A commonly used method for this is functional enrichment analysis, which identifies biological functions, pathways, and molecular activities that are overrepresented within a gene set when compared to a random selection of genes with a similar size and degree distribution from the genome. In our study, we performed Gene Ontology (GO) enrichment analysis using the R package clusterProfiler (Ashburner et al., 2000a), which leverages data from databases such as KEGG and GO to identify enriched biological terms. A greater degree of enrichment suggests that the network exhibits more meaningful gene interactions than would be expected by chance. This unique enrichment across subtypes implies that the gene networks represent biologically significant interactions, where genes within specific cancer subtype networks are functionally connected as a group. To evaluate the functional diversity between two gene networks, we conducted an experiment using GO to count the number of unique GO terms associated with the genes in each network. Specifically, we used the enrichGO() function from clusterProfiler to map the genes from both networks to their corresponding GO terms. The compareCluster() function was applied to compare the sets of GO terms associated with each network and to identify differences, focusing on the number of enriched biological functions. To quantify the differences, we calculated the number of enriched biological functions (#EBF) using the symmetric difference between the sets of GO terms. Mathematically, this is represented as: #$\#EBF=(GO(\mathcal{G}_{1})\setminus GO(\mathcal{G}_{2}))\cup(GO(\mathcal{G}_{2})\setminus GO(\mathcal{G}_{1}))$. This operation captures the unique functions present in one network but not another. Enrichment was evaluated based on statistical significance, where the biological functions with a p-value $<0.05$ were reported. A higher #EBF indicates that the networks capture different biological processes or molecular functions, potentially reflecting the underlying biological differences between the networks’ contexts.
Appendix G GO Function Enrichment Analysis
G.1 GO (Gene Ontology) and Biological Process
Gene Ontology (GO) (Ashburner et al., 2000a) is a comprehensive and standardized framework for annotating genes based on their functions in biological contexts. It provides a structured vocabulary to describe the roles of genes in three broad categories:
  • Biological Process (BP): This category defines the biological objectives or events that the gene products are involved in. BP terms focus on cellular functions that go awry in diseases like cancer. It includes essential processes such as “signal transduction”, “cell cycle regulation”, “immune response”, and “metabolic processes”.
  • Cellular Component (CC): This category defines the cellular locations where the gene products carry out their functions. CC terms focus on the spatial aspect of gene activity. Examples include terms like “cytosol”, “nucleus”, “mitochondrion”, and “plasma membrane”.
  • Molecular Function (MF): This category outlines the biochemical activities of the gene product. MF terms focus on how genes contribute to cellular machinery on a molecular level. Examples include terms like “ATP binding”, “enzyme activity”, and “protein binding”.
The Biological Process (BP) category is of particular relevance to cancer research because it directly reflects the underlying mechanisms that drive cancer development and progression and, therefore, serves as a suitable representative class for cancer-related GO terms (Desmedt et al., 2008). Identifying disruptions in BP terms related to cancer mechanisms can also help guide therapeutic strategies. For instance, drugs targeting cell cycle checkpoints, such as cyclin-dependent kinase inhibitors, or those promoting apoptosis, like Bcl-2 family inhibitors, are being developed to specifically correct or mitigate BP-related dysfunctions (Shapiro, 2006; Ashkenazi et al., 2017).
G.2 GO Function Enrichment Analysis Results
@@
Table 9 presents the enriched Gene Ontology (GO) terms associated with various biological functions across four cancer types (BRCA, GBM, LGG, and OV), as identified using different methods in the GO analysis of Experiment II. The Venn diagrams in Figure 3 illustrate the overlaps among the results from the different methods. Due to the complexity of comparing multiple methods, we present a four-way Venn diagram focusing on four selected methods (WGCNA, CSGNN, LR-GNN, and GeSubNet) for clarity.
GeSubNet findings: GeSubNet shows a robust array of enriched GO terms across different cancers, including:
@@
Comparison: The proposed method identifies a broader range of distinct GO terms compared to other methods, and the GO term set identified by GeSubNet constitutes a superset of the terms determined by different methods.
@@
Appendix H Simulated Gene Knockout Experiment
H.1 Workflow
@@ H.2 Shift Rate
Appendix I Case Study
I.1 Breast Invasive Carcinoma
@@
I.2 Experiments and Analysis Protocols
@@
Appendix J Ablation Studies
@@
Ablation. We conduct three detailed ablation studies to evaluate the impact of each module in GeSubNet. More details can be found in Appendix J. Figure 7 presents the results of the three ablation studies across all variant models. For Patient-M, the proposed sample encoder significantly outperforms all other DGM models across the four network inference tasks (BRCA, GBM, LGG, and OV). For instance, the proposed method achieves an average improvement of 32.3%/31.2%/22.1%/32.3% in terms of GED. The Graph-M ablations show that the method using Neo-GNN consistently performs best, while the other GNN models yield comparable results. For Infer-M ablation, GeSubNet significantly outperforms the other objective functions, achieving approximately twice the metric values of its counterparts.
@@
Appendix K Prior Graph V.S. Newly Generated Graph
We evaluated the performance of patient group learning by inputting the newly generated graph from the GeSubNet into a plain GCN and comparing the results. Figure 8 presents a UMAP visualization of the learned latent sample spaces, with the prior graph initialization (Left) and the generated graph GCN initialization (Right). The left sub-figure shows that different patient groups appear mixed in the latent sample space derived from the prior gene network. However, there are clearer boundaries between various patient groups, as shown on the right side. Such results confirm the redundancy of information in the common prior gene networks. It demonstrates that the GeSubNet provides more structured information and potential for cancer studies.